OpenStudy (anonymous):

which of the following estimates at 95% confidence level most likely comes from a small sample? a: 71%(-+4%) b:60%(-+18%) c:65%(-+2%) d:62%(-+6%)

OpenStudy (anonymous):

A confidence interval has the general form \[(\text{estimate}-\text{margin of error},~\text{estimate}+\text{margin of error})\] The margin of error is a value that depends on a constant (determined by the confidence level) and the standard error (which depends on the estimated statistic). For example, for a normally distributed population, a 95% confidence interval for the mean would be \[\left(\bar{x}-1.96\frac{\sigma}{\sqrt n},~\bar{x}+1.96\frac{\sigma}{\sqrt n}\right)\] where \(\bar{x}\) is the sample (estimate) mean, 1.96 is the critical value for a 95% confidence level (the constant I mentioned), \(\sigma\) is the standard deviation, and \(n\) is the sample size. Notice that if \(n\) is small, the fraction \(\dfrac{\sigma}{\sqrt n}\) tends to be large, whereas if \(n\) is large, the fraction approaches 0. From this you can gather that a smaller sample size generally increases the error margin.